OpenStudy (anonymous):

please show the work on how to solve x(x-8)=5?

OpenStudy (anonymous):

you can use quadratic equation

OpenStudy (anonymous):

\[x(x-8)=5 \Longrightarrow x^{2}-8x=5 \Longrightarrow x^{2}-8x-5=0\] then use quadratic formula

OpenStudy (anonymous):

I am not sure how to use the quadratic formula

OpenStudy (anonymous):

\[{-b \pm \sqrt{b^2-4ac}}\over{2a}\]

OpenStudy (anonymous):

ax^2+bx+c

OpenStudy (anonymous):

so this is my answer?

OpenStudy (anonymous):

no, this is not the final answer.... from the equation found by joemath, for your question a = 1, b = -8 and c = -5 also imranmeah has given the formula according to which - (-8) +/- root[(-8)^2 - 4(1)(-5) 8 +/- root[64 + 20] 8 +/- root(84) so x = ----------------------------- = ------------------- = --------------- 2 * 1 2 2 8 +/- 2root(21) 2 [4 +/- root(21)] x = --------------- = ----------------- 2 2 so two value of x are \[4 + \sqrt{21} and 4-\sqrt{21}\]